A solid is formed and it has 3 types of atoms X,Y,Z. X forms a FCC lattice with Y atoms occupying all the tetrahedral voids and Z atoms occupying half the octahedral voids. The formula of the solid is?

Grade:12

1 Answers

Ramreddy IIT Patna
9 years ago
no.of X atoms in unit cell= 4 (FCC)
no.of Y atoms in unit cell= 8 (All tetrahedral voids)
no.of X atoms in unit cell= 4 (All tetrahedral voids)
Formula of solid: XY2Z
….….….….….….….….….….….….….….….….….….….
….….….….….….….….….….….….….….….